What is the area of a square with diagonals 10cm long
Answer by checkley77(12844)

a^2+a^2=10^2
2a^2=100
a^2=100/2
a^2=50
a=sqrt50
a=7.07 cm is the length of the sides of the square.
7.07^2=50 cm^2